Circular No. 6301

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN 3 MPC 26211 contains observations of discrete components A, B, C and

possibly D by J. V. Scotti (Spacewatch) and A. Galád (Modra) during 1995 Dec. 27–29. Designated from west to east and with C the brightest, the components A, B and C can be associated with the three principal brightness peaks described by H. Böhnhardt and H. U. Käufl at the European Southern Observatory on Dec. 12 (IAUC 6274). Numerous further observations at seven additional observatories during 1995 Dec. 23–1996 Jan. 21 (notably a series by J. Chen and D. Jewitt with the University of Hawaii’s 2.2-m reflector during Dec. 23–27) are published on MPC 26444–26445. Böhnhardt and Häufl report that further observations with the New Technology Telescope show B separated from C by 4′′.6 (in p.a. 253^{o} on Jan. 7.0 UT and by 6′′.11 (in p.a. 244^{o}.8) on Jan. 31.0; the corresponding separations of A from C were 10′′.1 (in p.a. 247^{o}) and 17′′.05 (in p.a. 240^{o}.5). Furthermore, their deconvolution of images obtained by K. Reinsch and J. Storm with ESO 2.2-m and 1.5-m reflectors shows A at 1′′.4 (in p.a. 280^{o}) and 1′′.8 (in p.a. 270^{o}) from C on Nov. 28 and Dec. 2, respectively.

Z. Sekanina, Jet Propulsion Laboratory, communicates: “Solutions derived from the extensively tested model for split comets (Sekanina 1982, in Comets, University of Arizona Press, p. 251), applied to up to 13 of the most consistent observations obtained with large-aperture telescopes, indicate that the component B broke off from the primary component C most probably about 1995 Oct. 24 with a separation velocity of ∼ 1.3 m/s and with a relative nongravitational deceleration of 4.4 (in units of 10^{−5} of the solar attraction). This event was evidently followed by a secondary splitting of component B, which gave birth to component A on or about Dec. 1 (as independently suggested by Böhnhardt and Haüfl), with a separation velocity possibly as high as 4.6 m/s and a deceleration of, very crudely, 40 units. This B–A scenario offers a better fit to the observations than the alternative of having A separate from C in early November. There now appears to be no escape from the conclusion that the brightness outburst, which apparently occurred between Sept. 5 and 8 (IAUC 6227), preceded the first breakup episode by at least six weeks. As for component D, it seems that this might have separated from C in late November with an acceleration of more than 50 units. At least two further possible short-lived, high-acceleration condensations have been mentioned by other observers.”

1996 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden